Two angles of a triangle are in the ratio 3:4. The third angle is 15 degrees less than twice the first angle. Find all three angles.
Answer by ankor@dixie-net.com(22740)

Two angles of a triangle are in the ratio 3:4. The third angle is 15 degrees less than twice the first angle. Find all three angles.
:
Let:
1st angle = 3x
2nd angle = 4x
3rd angle = (2(3x) - 15)
3rd angle = (6x - 15)
:
3x + 4x + 6x - 15 = 180
13x = 180 + 15
13x = 195
x = 195/13
x = 15
:
Therefore our 3 angles;
3(15) = 45
4(15) = 60
6(15) - 15 = 75
:
:
Check: 45 + 60 + 75 = 180